# Mock Eisenstein series associated to partition ranks

### v1

@@ 2.4. The rank-crank PDE
The rank-crank PDE of Atkin and Garvan [3, Theorem 1.1] relates the rank and crank generating functions by a differential equation. To state it, we define the heat operator [^7] We point out that we re-normalize the standard heat operator by multiplying by a factor of $\frac{1}{4\pi^{2}}$ (see [11, p. 33] for more details). (of index $-\frac{3}{2}$) as
@@
Using the definition of $H$ and Lemma 2.1 with
$f(z;\tau)=\frac{\zeta^{\frac{1}{2}}(q)_{\infty}R(\zeta;q)}{1-\zeta}$
we obtain the lemma. ∎
@@ 2.5. Pólya cycle index polynomials
We require a result about Pólya cycle index polynomials in the case of symmetric group $S_{n}$, the set of permutations of the symbols $x_{1},x_{2},\ldots,x_{n}$; see [15] and [1, Lemma 2.1] for more details.
@@ 2.6. Divisibility of multinomials
For positive integers $a_{1},\ldots,a_{\ell}$ with sum $\sum_{j=1}^{\ell}a_{j}=n$, we have
@@
3. Mock Eisenstein series and the proof of Theorems 1.2 and 1.6
@@ 3.1. Modularity and completion
Motivated by Lemma 2.2, we define the mock Eisenstein series $f_{k}$ in terms of the rank generating function as follows:
@@
$R^{\circ}(z,\overline{z};\tau,\overline{\tau}):=$ $2i\sin(\pi z)q^{\frac{1}{24}}R^{\#}(z;\tau)e^{-12\pi^{2}G_{2}(\tau)z^{2}}$
=\ = $R(\zeta;q)+\frac{1}{2}q^{-\frac{1}{8}}\left(\zeta^{\frac{1}{2}}-\zeta^{-\frac{1}{2}}\right)\sum_{\pm}\pm\zeta^{\mp 1}S(3z\pm\tau;3\tau).$ (3.2)
@@
$\frac{c\tau+d}{2\pi iz}\exp\!\left(2\sum_{k,\ell\geq 0}f^{*}_{k,\ell}\left(\frac{a\tau+b}{c\tau+d},\frac{a\overline{\tau}+b}{c\overline{\tau}+d}\right)\frac{\left(\frac{2\pi iz}{c\tau+d}\right)^{k}}{k!}\frac{\left(\frac{2\pi i\overline{z}}{c\overline{\tau}+d}\right)^{\ell}}{\ell!}\right)$
@@ 3.2. Limiting behavior of $f_{k}$
Thus $S(3z\pm\tau;3\tau)$ is bounded as $\tau\to i\infty$. Multiplying with $(q)_{\infty}$ and taking the limit gives (3.7). The claim for $\widehat{f}_{k}$ follows from the claim for $f_{k}^{*}$ and Lemma 3.4. ∎
@@ 3.3. Differential equations
[…]
@@
Using Lemma 2.7 with $w=2\pi iz$ and ${x_{k}}=\frac{2}{k!}(3G_{k}-f_{k})$ we obtain
@@
$\frac{2}{(2\pi iz)^{3}}\exp\!\left(6\sum_{k\geq 1}G_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right)\\ =\left(6q\frac{\partial}{\partial q}+\left(\zeta\frac{\partial}{\partial\zeta}\right)^{2}+6G_{2}(\tau)\right)\frac{1}{2\pi iz}\exp\!\left(2\sum_{k,\ell\geq 0}f^{*}_{k,\ell}(\tau,\overline{\tau})\frac{(2\pi iz)^{k}}{k!}\frac{(2\pi i\overline{z})^{\ell}}{\ell!}\right).$
@@
$\frac{2}{(2\pi iz)^{3}}\exp\!\left(\!6\sum_{k\geq 1}G_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\!\right)\\ =\left(\!6q\frac{\partial}{\partial q}+\left(\zeta\frac{\partial}{\partial\zeta}\right)^{2}+6G_{2}(\tau)\!\right)\!\frac{1}{2\pi iz}\exp\!\left(\!2\sum_{k\geq 0}f^{*}_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\!\right)\!.$
@@
plugging in $y=\frac{z-\overline{z}}{2i}$. Using (3.11) and (3.6), we obtain
@@
using (3.6). The sum on the right-hand side equals $-\eta(-\overline{\tau})$, so the above becomes
@@
(4) The Serre derivative $\vartheta_{k}$ defined in (2.4) increases the weight of a modular object by $2$ (note that the proof in [17, p. 48] goes through if $f$ is non-holomorphic). We have
@@
$R(\zeta;q)=\frac{\sin\!\left(\pi z\right)}{\pi z(q)_{\infty}}\sum_{k\geq 0}\mathrm{Tr}_{k}(\phi,f;\tau)(2\pi iz)^{k}.$
Expanding $R(\zeta;q)=\sum_{k\geq 0}R_{k}(q)\frac{(2\pi iz)^{k}}{k!}$, letting $z\mapsto\frac{z}{2\pi i}$, and using that $\sin(-\frac{iz}{2})=-i\sinh(\frac{z}{2})$, we conclude
@@
[…]
@@
Both sides of the lemma are 00 for $k$ odd. Namely, since $N(m,n)=N(-m,n)$, we have that $R_{k}$ is 00 for $k$ odd, and for $\ell$ odd, we have
@@
$\!=\!\sum_{m\geq 3n\geq 3}(2m-6n)^{k}\!\left(\!-q^{n(2m-1)}+q^{(2n-1)(m-2)}+q^{n(2m+1)}-q^{(2n-1)(m-1)}\!\right)\!,$
making the change of variables $m\mapsto m-3n$. Interchanging the role of $m$ and $n$ in the second and fourth sum the above becomes
$-\sum_{m\geq 3n\geq 3}(2m-6n)^{k}q^{n(2m-1)}+\sum_{n\geq 3m\geq 3}(2n-6m)^{k}q^{(2m-1)(n-2)}\\ +\sum_{m\geq 3n\geq 3}(2m-6n)^{k}q^{n(2m+1)}-\sum_{n\geq 3m\geq 3}(2n-6m)^{k}q^{(2m-1)(n-1)}.$
@@
$-\sum_{2m-1\geq 6n-1\geq 5}(2m-1-6n+1)^{k}q^{n(2m-1)}+\sum_{2n+1\geq 3(2m-1)\geq 3}(2n+1-3(2m-1))^{k}q^{n(2m-1)}$
$+\sum_{2m-1\geq 6n+1\geq 7}(2m-1-6n-1)^{k}q^{n(2m-1)}-\sum_{2n-1\geq 3(2m-1)\geq 3}(2n-1-3(2m-1))^{k}q^{n(2m-1)}$
$\qquad=\sum_{\begin{subarray}{c}m\geq 6n+1\geq 7\\ \text{$m$ odd}\end{subarray}}\left((m-6n-1)^{k}-(m-6n+1)^{k}\right)q^{nm}$
$\hskip 142.26378pt+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\\ \text{$m$ odd}\end{subarray}}\left((2n-3m+1)^{k}-(2n-3m-1)^{k}\right)q^{nm}$
$\qquad=\sum_{\begin{subarray}{c}m\geq 6n+1\geq 7\end{subarray}}\left((m-6n-1)^{k}-(m-6n+1)^{k}\right)q^{nm}$
$\hskip 142.26378pt+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\end{subarray}}\left((2n-3m+1)^{k}-(2n-3m-1)^{k}\right)q^{nm}.$ (4.3)
@@
[…]
@@
Finally, we refine Lemma 3.5 by computing the first Fourier coefficients of $f_{k}$. In particular, the $f_{k}$ are naturally normalized in the sense that the second Fourier coefficient is $1$.
@@ 5. Examples
Here, we write down the first Fourier coefficients of the mock Eisenstein series $f_{k}$. We have
@@ 6. Questions for future research
    The three properties of the $f_{k}$ given in Theorem 1.2 do not determine them uniquely. Therefore, we provide two recursive definitions of $f_{k}$ in Theorem 1.4. It would be interesting to find “nice” properties that define a mock Eisenstein series uniquely. For example, the fact that $G_{k}$ is a normalized Hecke eigenform of weight $k$ which does not vanish at $i\infty$ determines it uniquely. As the coefficient of $q$ in the Fourier expansion of the $f_{k}$ vanishes by Proposition 4.3, one deduces that $f_{k}$ cannot be a Hecke eigenform.
@@
    The functions $f_{2},f_{4},f_{6},\ldots$ together with $G_{2},G_{4}$ and $G_{6}$, do not seem to satisfy any algebraic relations. This has been verified numerically up to weight 24 (and also up to mixed weight 12). Is it indeed the case that the algebra $\mathcal{F}$ is free?
@@
    $\sum_{\begin{subarray}{c}an-1\geq bm\geq b\end{subarray}}(an-bm)^{\ell-1}q^{mn}-\sum_{\begin{subarray}{c}n-1\geq abm\geq ab\end{subarray}}(n-abm)^{\ell-1}q^{mn}$
    a mock Eisenstein series? Note that $a=2$ and $b=3$ yields the function $g_{\ell}$ in this paper. Are other choices of $a$ and $b$ also of particular interest? Do these functions, together with those in [18, p.15] and [13, equation (1.4)], form the first examples of a theory of (higher level) mock Eisenstein series?

### v2

@@ 2.4. The rank-crank PDE
The rank-crank PDE of Atkin and Garvan [3, Theorem 1.1] relates the rank and crank generating functions by a differential equation. To state it, we define the heat operator [^8] We point out that we re-normalize the standard heat operator by multiplying by a factor of $\frac{1}{4\pi^{2}}$ (see [14, p. 33] for more details). (of index $\frac{3}{2}$) as
@@
We obtain the lemma, using the definition of $H$ and Lemma 2.1 with
$f(z;\tau)=\frac{\zeta^{\frac{1}{2}}(q)_{\infty}R(\zeta;q)}{1-\zeta}.\qed$
@@ 2.5. Pólya cycle index polynomials
We require a result about Pólya cycle index polynomials in the case of symmetric group $S_{n}$, the set of permutations of the symbols $x_{1},x_{2},\ldots,x_{n}$, see [22] and [1, Lemma 2.1] for more details.
@@ 2.6. Divisibility of multinomials
For $a_{1},\ldots,a_{\ell}\in\mathbb{N}$ with $\sum_{j=1}^{\ell}a_{j}=n$, we have
@@
3. Eisenstein–type series and the proof of Theorems 1.2 and 1.6
@@ 3.1. Modularity and completion
Motivated by Lemma 2.2, we define the Eisenstein–type series $f_{k}$ in terms of the rank generating function as follows:
@@
$R^{\circ}(z,\overline{z};\tau,\overline{\tau}):=\$$2i\sin(\pi z)q^{\frac{1}{24}}R^{\#}(z;\tau)e^{-12\pi^{2}G_{2}(\tau)z^{2}}$
[…]
@@
$\frac{c\tau+d}{2\pi iz}\exp\!\left(2\sum_{k,\ell\geq 0}f^{*}_{k,\ell}\left(\frac{a\tau+b}{c\tau+d},\frac{a\overline{\tau}+b}{c\overline{\tau}+d}\right)\frac{\left(\frac{2\pi iz}{c\tau+d}\right)^{k}}{k!}\frac{\left(\frac{2\pi i\overline{z}}{c\overline{\tau}+d}\right)^{\ell}}{\ell!}\right)\hskip-136.5733pt$
@@ 3.2. Limiting behavior of $f_{k}$
Thus $q^{-\frac{1}{8}}S(3z\pm\tau;3\tau)\to 0$ as $\tau\to i\infty$, from which we obtain (3.7). The claim for $\widehat{f}_{k}$ follows from the claim for $f_{k}^{*}$ and Lemma 3.4. ∎
@@ 3.3. Differential equations
(1) It is well-known, and follows by (2.3), that $D(G_{\ell})\in\mathcal{F}$. Thus it remains to show the formula for $D(f_{k})$ for $k\in 2\mathbb{N}$ which directly implies that $D(f_{k})\in\mathcal{F}$. Using Theorem 2.5, Lemma 2.2, and (3.1), we obtain
@@
Using Lemma 2.7 with $w=2\pi iz$ and ${x_{k}}=\frac{2}{k!}(3G_{k}-f_{k})$, we obtain
@@
$\frac{2}{(2\pi iz)^{3}}\exp\!\left(6\sum_{k\geq 1}G_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\right)\\ =\left(6q\frac{\partial}{\partial q}+\left(\zeta\frac{\partial}{\partial\zeta}\right)^{2}+6G_{2}(\tau)\right)\left(\frac{1}{2\pi iz}\exp\!\left(2\sum_{k,\ell\geq 0}f^{*}_{k,\ell}(\tau,\overline{\tau})\frac{(2\pi iz)^{k}}{k!}\frac{(2\pi i\overline{z})^{\ell}}{\ell!}\right)\right).$
@@
$\frac{2}{(2\pi iz)^{3}}\exp\!\left(\!6\sum_{k\geq 1}G_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\!\right)\\ =\left(\!6q\frac{\partial}{\partial q}+\left(\zeta\frac{\partial}{\partial\zeta}\right)^{2}+6G_{2}(\tau)\!\right)\!\left(\frac{1}{2\pi iz}\exp\!\left(\!2\sum_{k\geq 0}f^{*}_{k}(\tau)\frac{(2\pi iz)^{k}}{k!}\!\right)\right)\!.$
@@
using that $y=\frac{z-\overline{z}}{2i}$. Using (3.11) and (3.6), we obtain
@@
using (3.6). The sum on the right-hand side equals $\eta(-\overline{\tau})$, so the above becomes
@@
(4) The Serre derivative $\vartheta_{k}$ defined in (2.4) increases the weight of a modular object by $2$ (note that the proof in [24, p. 48] goes through if $f$ is non-holomorphic). We have
@@
$R(\zeta;q)=\frac{\sin\!\left(\pi z\right)}{\pi z(q)_{\infty}}\sum_{k\geq 0}\mathrm{Tr}_{k}(\phi,f;\tau)(2\pi iz)^{k}.$ (3.13)
From (1.1) and (1.5) we deduce that
$R(\zeta;q)=\sum_{k\geq 0}R_{k}(q)\frac{(2\pi iz)^{k}}{k!}.$ (3.14)
Substituting $z\mapsto\frac{z}{2\pi i}$ and applying (3.14) and $\sin(-\frac{iz}{2})=-i\sinh(\frac{z}{2})$ in (3.13), we obtain
@@
$R_{k}(q)=\frac{2^{2-k}}{(q)_{\infty}}\sum_{\begin{subarray}{c}\ell=2\\ \ell\equiv k\ \left(\mathrm{mod}\,2\right)\end{subarray}}^{k}\binom{k}{\ell-1}\left(g_{\ell}(\tau)+\left(2^{\ell-1}-1\right)\frac{B_{\ell}}{2\ell}\right).$
@@
Both sides of the lemma are zero for $k$ odd. Namely, since N(m,n)=N(m,n)N(m,n)=N(-m,n), we have that $R_{k}$ is zero for $k$ odd, and for $\ell$ odd, we have
@@
$\!=\!\hskip-4.26773pt\sum_{m\geq 3n\geq 3}\hskip-4.26773pt(2m-6n)^{k}\!\left(\!-q^{n(2m-1)}+q^{(2n-1)(m-2)}+q^{n(2m+1)}-q^{(2n-1)(m-1)}\!\right)\!,$
making the change of variables $m\mapsto m-3n$. Interchanging the role of $n$ and $m$ in the second and fourth sum the above becomes
$\sum_{m\geq 3n\geq 3}(2m-6n)^{k}q^{n(2m-1)}+\sum_{n\geq 3m\geq 3}(2n-6m)^{k}q^{(n-2)(2m-1)}\\ +\sum_{m\geq 3n\geq 3}(2m-6n)^{k}q^{n(2m+1)}-\sum_{n\geq 3m\geq 3}(2n-6m)^{k}q^{(n-1)(2m-1)}.$
@@
$-\hskip-8.5359pt\sum_{2m-1\geq 6n-1\geq 5}\hskip-8.5359pt(2m-1-6n+1)^{k}q^{n(2m-1)}+\hskip-8.5359pt\sum_{2n+1\geq 3(2m-1)\geq 3}\hskip-8.5359pt(2n+1-3(2m-1))^{k}q^{n(2m-1)}$
$+\hskip-8.5359pt\sum_{2m-1\geq 6n+1\geq 7}\hskip-8.5359pt(2m-1-6n-1)^{k}q^{n(2m-1)}-\hskip-8.5359pt\sum_{2n-1\geq 3(2m-1)\geq 3}\hskip-8.5359pt(2n-1-3(2m-1))^{k}q^{n(2m-1)}$
$\qquad=\sum_{\begin{subarray}{c}m\geq 6n+1\geq 7\\ \text{$m$ odd}\end{subarray}}\hskip-2.84544pt\left((m-6n-1)^{k}-(m-6n+1)^{k}\right)q^{nm}$
$\hskip 142.26378pt+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\\ \text{$m$ odd}\end{subarray}}\hskip-2.84544pt\left((2n-3m+1)^{k}-(2n-3m-1)^{k}\right)q^{nm}$
$\qquad=\sum_{\begin{subarray}{c}m\geq 6n+1\geq 7\end{subarray}}\hskip-2.84544pt\left((m-6n-1)^{k}-(m-6n+1)^{k}\right)q^{nm}$
$\hskip 142.26378pt+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\end{subarray}}\hskip-2.84544pt\left((2n-3m+1)^{k}-(2n-3m-1)^{k}\right)q^{nm}.$ (4.3)
@@
$\hskip 28.45274pt=\sum_{{n\geq 0}}B_{n}\!\left(\frac{1}{2}\right)\frac{(2\pi iz)^{n}}{n!}\sum_{k\geq 1}2^{2-k}\sum_{\begin{subarray}{c}\ell=1\\ \ell\not\equiv k\ \,(\mathrm{mod}\,2)\end{subarray}}^{k}\binom{k}{\ell}\left(g_{\ell+1}(\tau)+\frac{\left(2^{\ell}-1\right)B_{\ell+1}}{2(\ell+1)}\right)\frac{(2\pi iz)^{k}}{k!}$
$\hskip 28.45274pt=\sum_{\begin{subarray}{c}n\geq 0\\ \ell\geq 1\end{subarray}}\sum_{\begin{subarray}{c}k\geq\ell\\ k\not\equiv\ell\ \,(\mathrm{mod}\,2)\end{subarray}}\binom{n}{k}\binom{k}{\ell}B_{n-k}\!\left(\frac{1}{2}\right)2^{2-k}\left(g_{\ell+1}(\tau)+\frac{\left(2^{\ell}-1\right)B_{\ell+1}}{2(\ell+1)}\right)\frac{(2\pi iz)^{n}}{n!},$
@@
Finally, we refine Lemma 3.5 by computing the first Fourier coefficients of $f_{k}$. In particular, the $f_{k}$ are naturally normalized in the sense that the second Fourier coefficient is one.
@@
Remark.
Using a similar argument as in the proof of Propostion 4.3, one can show that the $n$-th Fourier coefficient of $f_{k}$ is a linear combination of the $k$-th powers of $\{1,\ldots,n-1\}$.
@@ 5. Examples
Here, we write down the first Fourier coefficients of the Eisenstein–type series $f_{k}$. We have
@@ 6. Questions for future research
    The three properties of the $f_{k}$ given in Theorem 1.2 do not determine them uniquely. Therefore, we provide two recursive definitions of $f_{k}$ in Theorem 1.3. It would be interesting to find “nice” properties that define an Eisenstein–type series uniquely. For example, the fact that $G_{k}$ is a normalized Hecke eigenform of weight $k$ which does not vanish at $i\infty$ determines it uniquely. As the coefficient of $q$ in the Fourier expansion of the $f_{k}$ vanishes by Proposition 4.3, one deduces that $f_{k}$ cannot be a Hecke eigenform with respect to the usual action of the Hecke operators on $q$-series.
@@
    The functions $f_{2},f_{4},f_{6},\ldots$ together with $G_{2},G_{4}$, and $G_{6}$, do not seem to satisfy any algebraic relations. This has been verified numerically up to weight 24 (and also up to mixed weight 12). Is it indeed the case that the algebra $\mathcal{F}$ is free?
@@
    $\sum_{\begin{subarray}{c}an-1\geq bm\geq b\end{subarray}}(an-bm)^{\ell-1}q^{nm}-\sum_{\begin{subarray}{c}n-1\geq abm\geq ab\end{subarray}}(n-abm)^{\ell-1}q^{nm}$
    an Eisenstein–type series? Note that $a=2$ and $b=3$ yields the function $g_{\ell}$ in this paper. Are other choices of $a$ and $b$ also of particular interest? Do these functions, together with those in [25, p.15] and [19, equation (1.4)], form the first examples of a theory of (higher level) Eisenstein–type series?